# Self-Adaptive Graph Mixture of Models

### v1

@@ 1 Introduction
Graph Neural Network (GNN) architectures have seen substantial progress since their early formulation in seminal work on graph-based semi-supervised learning (kipf2016semi). Though developments in the neural architecture have led to an improved performance (on an average), there is still no consensus on distinct advantage of one model over others. This is reflected in two important observations in recent times: (a) the performance of GNN models have plateaued with no significant performance gain reported with complex architecture. (b) slight tuning of hyperparameters of classical GNN models like GCN (kipf2016semi), GAT (velivckovic2017graph), and GraphSAGE (hamilton2017inductive) have shown state-of-the-art (SOTA) performance in node classification tasks, matching or even surpassing latest Graph Transformers (GTs) across diverse datasets (luo2024classic). This suggests different models tend to learn different (overlapping) regions of the representation space, but each falling short of covering the space to capture diverse patterns and features of the datasets. Moreover, selecting the right model for a dataset often involves trial-and-error and computationally expensive hyperparameter tuning, with many models discarded after underperforming. However, each model may still capture unique structural patterns or inductive biases (model assumptions) that, while insufficient alone, could contribute meaningfully when combined. In machine learning (ML), ensemble learning approach is a proven method to pool a set of (weak) models to create a strong model (li2021neural). In recent years, mixture of experts (MoE) approach has become the de facto choice for ensemble learning, though it differs from the classical approach by the gating (routing) network which is trained to select expert(s) based on the input context. The MoE and its variants have been successfully applied to large language models (LLMs) to increase model capacity without proportionate increase in computational cost (fedus2022switch). In the graph domain, MoE-based approaches are still in the early stages of exploration but have demonstrated significant potential. Existing methods such as GMoE (wang2023graphmixtureexpertslearning) typically exhibit limited architectural diversity, often employing variations of a single base model (e.g., GCNs with MoE layers inserted at intermediate depths). Furthermore, their gating mechanisms are generally not designed to capture complex graph topologies, and expert pruning strategies, if present, are either ad hoc or inefficiently integrated. These limitations underscore the need for a more flexible and topology-aware MoE framework tailored to the unique challenges of graph-structured data.
@@
In this work, we propose Self-Adaptive Graph Mixture of Models (SAGMM), a modular and practical framework that learns to automatically select and combine the most appropriate GNN models for each part of the graph. SAGMM leverages a diverse pool of GNN architectures, each serving as an expert with a fixed view of the data, and employs a topology-aware sparse attention gate that dynamically routes input to the most relevant experts. The framework is designed to be trained end-to-end but as a practical variation, we also explore a training strategy, where experts are pre-trained independently and only the routing mechanism and task head are trained, reducing training time. The core idea of SAGMM is that different GNNs offer complementary perspectives of the graph, but their architectural rigidity limits their individual expressiveness. By selectively combining them at the representation level, SAGMM achieves greater adaptability and structural awareness. This formulation also aligns with the broader model merging paradigm (yang2024model), though our merging occurs in representation space rather than parameter space. The salient features of the SAGMM are as follows:
@@ 2.2 Graph Neural Networks
Several GNN architectures based on encoder-decoder framework (hamilton2017representation) with encoding performed by message-passing mechanism have been proposed for graph ML tasks. According to (balcilar2021analyzing), the general update rule for the node embeddings $H^{(l+1)}$ at layer $l+1$ is represented by:
@@
where $\sigma(\cdot)$ is a nonlinear activation function, $W^{(l,q)}$ are learnable parameters associated with the filter $q$, and $C^{(l,q)}$ are model-specific convolution matrices. Different GNNs define $C^{(l,q)}$ differently. For instance, GCN (kipf2016semi) uses a normalized adjacency matrix fixed across layers, GraphSAGE (hamilton2017inductive) uses identity and mean aggregation, GAT (velivckovic2017graph) introduces attention-based weighting, while GIN (xu2018powerful) approximates the Weisfeiler-Lehman (WL) test via learnable aggregation (ding2022sketchgnn). JKNet (xu2018representation) improves depth flexibility by aggregating information from multiple GNN layers via jump connections. Variants such as SGC (wu2019simplifyinggraphconvolutionalnetworks) eliminate nonlinearities for efficiency, MixHop (abu2019mixhop) aggregates multi-hop neighborhoods, and GraphCNN (defferrard2016convolutional) applies spectral filtering via Chebyshev polynomials.
@@ 2.3 Mixture of Experts in Graph Learning
The MoE framework (jacobs1991adaptive), when first introduced enables routing inputs to specialized models via a trainable dense gating network. Modern sparse activation MoE architectures, widely used in large-scale language and vision models, leverage thousands of experts to enhance model capacity (zhang2025mixture). Methods such as DynMoE (guo2024dynamic) introduces top-any gating and pruning strategy to dynamically activate and prunes expert in LLM domain. While M32 (wu2024yuan) introduces an attention-based routing mechanism, its quadratic complexity limits scalability for large graphs. Recent surveys (cai2025survey; zhang2025mixture) provide a comprehensive overview of trends and challenges in MoE models for LLMs.
For graph data, GMoE (wang2023graphmixtureexpertslearning) extends MoE to GNNs, enabling multi-hop information aggregation. GraphMETRO (wu2024graphmetro) leverages MoE to tackle distribution shifts in GNNs. TopExpert (kim2023learningtopologyspecificexpertsmolecular) employs domain-specific linear experts for graph classification but lacks adaptability to diverse tasks. DA-MoE (yao2024moe) uses GNN layers as experts to address depth sensitivity but does not support dynamic expert selection. GraphMoRE (guo2025graphmore) leverages a Mixture of Riemannian Experts, where each expert operates on a learned manifold to capture diverse local structures. Link-MoE (ma2024mixture) relies on pre-trained GNN experts for link prediction, limiting its flexibility for broader applications.
@@ 3 Methodology
We now introduce the SAGMM in detail, with the subsections focusing its key components. Figure 1 illustrates the SAGMM architecture. Input features are first enhanced with structural context, and expert selection is performed by a gating network using Simple Global Attention (SGA) (wu2023sgformer) and learnable thresholds. The selected expert outputs are projected into a shared space, weighted by gate scores, and aggregated, while low-importance experts are pruned based on an adaptive threshold $\eta$. The training phase of SAGMM in outlined in Algorithm 1.
@@ 3.1 Pool of Experts
Recent surveys (chen2020bridging; ju2024comprehensive) categorize GNNs along three key dimensions: propagation strategy (spectral vs. spatial), aggregation mechanism (mean vs. attention), and training setup (transductive vs. inductive). Guided by these dimensions, we include GNNs like GCN (kipf2016semi) for its spectral filtering and suitability for homophilic graphs, GAT (velivckovic2017graph) for attention-based neighbor weighting, and GraphSAGE (hamilton2017inductive) for its inductive capability through neighborhood sampling, along with additional models in the expert pool. Such heterogeneity enables the router to select experts that align with the structural and statistical properties of individual instances.
@@
This design is also motivated by the bias-variance tradeoff (yang2020rethinking), where over-parameterized models can improve generalization by reducing variance. By incorporating multiple GNNs, our framework benefits from this property. Moreover, in line with the No Free Lunch theorem (goldblum2023no) which states that no single model performs optimally across all tasks, our approach accounts for variations in network topology, homophily, and heterophily that affects GNN performance. We show empirically that by pooling experts, each with an unique message-passing mechanism, within MoE framework, our approach dynamically adapts to diverse graph structures, overcoming limitations of a single model.
@@ 3.2 Gating Network
Limitations of traditional gating mechanism. Several approaches have been proposed for designing gating networks in MoE architectures (jordan1994hierarchical; clark2022unified; zhou2202mixture). A widely adopted sparse gating strategy is the noisy top-$k$ gating mechanism (shazeer2017outrageously), which applies a softmax over the top-$k$ logits produced by a linear transformation of the input. To encourage expert diversity and prevent load imbalance, this approach injects learnable Gaussian noise into the logits, a modification later adapted for graph data in GMoE model.
@@
    The model’s performance is highly dependent on the choice of $k$, as different datasets exhibit varying optimal values as shown in Figure 2(a) for GMoE-GCN (wang2023graphmixtureexpertslearning), highlighting the need for extensive hyperparameter tuning. Moreover, using a fixed $k$ forces all nodes to activate the same number of experts, which may not be needed considering structural and feature characteristics. As shown in Figure 2(b), different nodes activate different number of experts as per their requirement, eventually leading to performance gains.
@@
    Existing methods often rely on simple linear projections of node features to compute gating scores, neglecting structural dependencies within the graph. While DA-MoE (yao2024moe) and GraphMoRE (guo2025graphmore) attempt to overcome this limitation by leveraging GNN-based gating networks, a fundamental drawback remains: these approaches are still unable to effectively capture global graph structural information, which is vital for informed expert selection.
To address these limitations, for each node, TAAG selects experts based on the local and global structural information while dynamically adjusting the number of activated experts to improve efficiency and robustness. We leverage the SGA block from SGFormer (wu2023sgformer) to compute attention scores efficiently. While standard attention mechanisms have a quadratic complexity of $O(n^{2})$, SGA operates in linear time $O(n)$ w.r.t. $n$ nodes, making it scalable for medium to large-scale graphs. To incorporate global structural information, we construct the matrix $X^{(g)}$ using the $p$ smallest eigenvectors of the normalized graph Laplacian $L$ (Eq. (4)), providing positional encodings for each node (rampasek2022GPS). We do not use diffusion-based methods (gasteiger2019diffusion; chamberlain2021grand) as they involve expensive preprocessing, assumes homophily and perform poorly in tasks like link prediction. To further strengthen the structural information, we also add local structural features by aggregating 1-hop and 2-hop (approximate) features (hamilton2017inductive; wang2020multi). The detailed formulations are provided in Eq. (3) and Eq. (4).
@@
where: $W_{Q},W_{K},W_{V}\in\mathbb{R}^{(d+p)\times N}$ are learnable projection matrices for query, key, and value, respectively. We compute the attention-based gating scores using the Simple Global Attention (SGA) mechanism (wu2023sgformer):
@@ 3.3 Adaptive Expert Pruning
Auxiliary Loss Functions. To ensure balanced expert utilization, we incorporate importance loss and diversity loss from prior works (wang2023graphmixtureexpertslearning; guo2024dynamic). These losses prevent over-reliance on specific experts by encouraging more even and orthogonal expert activation patterns.
@@ 4 Experiments
Initial Number of Experts ($N_{0}$). For our primary task of node classification, we construct the expert pool by selecting four and eight widely recognized GNN architectures: GCN, JKNet, GraphCNN, MixHop, GAT, SGC, GIN, and GraphSAGE, all of which are established standards for this application (hu2020open). In the case of other tasks, the expert pool is composed of four GNN models. The decision to use either eight or four experts is guided by the need to balance architectural diversity with computational feasibility. Notably, this choice is consistent with recent findings in the literature (yun2024toward; he2025efficiently), which indicate that configuring MoE models with four or eight experts enables efficient inference and strong model performance.
@@ 4.1 Performance Evaluation
Node Classification. We evaluate on seven standard datasets ranging from small to large-scale, including four homophilic graphs: ogbn-arxiv, ogbn-proteins, ogbn-products, and ogbn-papers100M from the OGB benchmark (hu2020open), and three heterophilic graphs: Pokec (jure2014snap), Deezer (rozemberczki2020characteristic), and YelpChi (mukherjee2013yelp). As shown in Table 1, SAGMM outperforms all baselines, achieving notable gains on medium to large-scale datasets such as a 5.90% improvement on ogbn-products and 1.57% on ogbn-papers100M, along with consistent accuracy improvements across the remaining benchmarks. Our analysis reveals that SAGMM consistently prunes weaker experts such as GIN across datasets, demonstrating its ability to adaptively select the most effective experts.
@@ 4.2 Ablation Study
We conduct an ablation study to evaluate the contribution of each component in SAGMM by testing four variants: (i) using identical GNN architectures for all experts (SAGMM (w/o diverse)), (ii) replacing our gating module with noisy Top-$k$ gating (SAGMM (top-$k$ gating)), (iii) using Top-Any gating (guo2024dynamic) having dynamic expert selection (SAGMM (top-any gating)), and (iv) disabling the adaptive expert pruning mechanism (SAGMM (w/o pruning)). As shown in Table 5, the largest drop is observed when expert diversity is removed, confirming its critical role. Substituting our gating mechanism with top-$k$ or top-any variants impairs expert selection quality, while disabling pruning retains performance in the case of ogbl-collab but always leads to higher memory usage. These results underscore the importance of expert diversity, TAAG gating, and pruning in SAGMM’s effectiveness.

### v2

@@ 1 Introduction
Graph Neural Network (GNN) architectures have seen substantial progress since their early formulation in seminal work on graph-based semi-supervised learning (Kipf and Welling 2016). Though developments in the neural architecture have led to an improved performance (on an average), there is still no consensus on distinct advantage of one model over others. This is reflected in two important observations in recent times: (a) the performance of GNN models have plateaued with no significant performance gain reported with complex architecture. (b) slight tuning of hyperparameters of classical GNN models like GCN (Kipf and Welling 2016), GAT (Veličković et al. 2017), and GraphSAGE (Hamilton et al. 2017a) have shown state-of-the-art (SOTA) performance in node classification tasks, matching or even surpassing latest Graph Transformers (GTs) across diverse datasets (Luo et al. 2024). This suggests different models tend to learn different (overlapping) regions of the representation space, but each falling short of covering the space to capture diverse patterns and features of the datasets. Moreover, selecting the right model for a dataset often involves trial-and-error and computationally expensive hyperparameter tuning, with many models discarded after underperforming. However, each model may still capture unique structural patterns or inductive biases (model assumptions) that, while insufficient alone, could contribute meaningfully when combined. In machine learning (ML), ensemble learning approach is a proven method to pool a set of (weak) models to create a strong model (Li et al. 2021). In recent years, mixture of experts (MoE) approach has become the de facto choice for ensemble learning, though it differs from the classical approach by the gating (routing) network which is trained to select expert(s) based on the input context. The MoE and its variants have been successfully applied to large language models (LLMs) to increase model capacity without proportionate increase in computational cost (Fedus et al. 2022). In the graph domain, MoE-based approaches are still in the early stages of exploration but have demonstrated significant potential. Existing methods such as GMoE (Wang et al. 2023) typically exhibit limited architectural diversity, often employing variations of a single base model (e.g., GCNs with MoE layers inserted at intermediate depths). Furthermore, their gating mechanisms are generally not designed to capture complex graph topologies, and expert pruning strategies, if present, are either ad hoc or inefficiently integrated. These limitations underscore the need for a more flexible and topology-aware MoE framework tailored to the unique challenges of graph-structured data.
@@
In this work, we propose Self-Adaptive Graph Mixture of Models (SAGMM), a modular and practical framework that learns to automatically select and combine the most appropriate GNN models for each part of the graph. SAGMM leverages a diverse pool of GNN architectures, each serving as an expert with a fixed view of the data, and employs a topology-aware sparse attention gate that dynamically routes input to the most relevant experts. The framework is designed to be trained end-to-end but as a practical variation, we also explore a training strategy, where experts are pre-trained independently and only the routing mechanism and task head are trained, reducing training time. The core idea of SAGMM is that different GNNs offer complementary perspectives of the graph, but their architectural rigidity limits their individual expressiveness. By selectively combining them at the representation level, SAGMM achieves greater adaptability and structural awareness. This formulation also aligns with the broader model merging paradigm (Yang et al. 2024), though our merging occurs in representation space rather than parameter space. The salient features of the SAGMM are as follows:
@@ 2.2 Graph Neural Networks
Several GNN architectures based on encoder-decoder framework (Hamilton et al. 2017b) with encoding performed by message-passing mechanism have been proposed for graph ML tasks. According to (Balcilar et al. 2021), the general update rule for the node embeddings $H^{(l+1)}$ at layer $l+1$ is represented by:
@@
where $\sigma(\cdot)$ is a nonlinear activation function, $W^{(l,q)}$ are learnable parameters associated with the filter $q$, and $C^{(l,q)}$ are model-specific convolution matrices. Different GNNs define $C^{(l,q)}$ differently. For instance, GCN (Kipf and Welling 2016) uses a normalized adjacency matrix fixed across layers, GraphSAGE (Hamilton et al. 2017a) uses identity and mean aggregation, GAT (Veličković et al. 2017) introduces attention-based weighting, while GIN (Xu et al. 2018a) approximates the Weisfeiler-Lehman (WL) test via learnable aggregation (Ding et al. 2022). JKNet (Xu et al. 2018b) improves depth flexibility by aggregating information from multiple GNN layers via jump connections. Variants such as SGC (Wu et al. 2019) eliminate nonlinearities for efficiency, MixHop (Abu-El-Haija et al. 2019) aggregates multi-hop neighborhoods, and GraphCNN (Defferrard et al. 2016) applies spectral filtering via Chebyshev polynomials.
@@ 2.3 Mixture of Experts in Graph Learning
The MoE framework (Jacobs et al. 1991), when first introduced enables routing inputs to specialized models via a trainable dense gating network. Modern sparse activation MoE architectures, widely used in large-scale language and vision models, leverage thousands of experts to enhance model capacity (Zhang et al. 2025). Methods such as DynMoE (Guo et al. 2024) introduces top-any gating and pruning strategy to dynamically activate and prunes expert in LLM domain. While M32 (Wu et al. 2024a) introduces an attention-based routing mechanism, its quadratic complexity limits scalability for large graphs. Recent surveys (Cai et al. 2025; Zhang et al. 2025) provide a comprehensive overview of trends and challenges in MoE models for LLMs.
For graph data, GMoE (Wang et al. 2023) extends MoE to GNNs, enabling multi-hop information aggregation. GraphMETRO (Wu et al. 2024b) leverages MoE to tackle distribution shifts in GNNs. TopExpert (Kim et al. 2023) employs domain-specific linear experts for graph classification but lacks adaptability to diverse tasks. DA-MoE (Yao et al. 2024) uses GNN layers as experts to address depth sensitivity but does not support dynamic expert selection. GraphMoRE (Guo et al. 2025) leverages a Mixture of Riemannian Experts, where each expert operates on a learned manifold to capture diverse local structures. Link-MoE (Ma et al. 2024) relies on pre-trained GNN experts for link prediction, limiting its flexibility for broader applications.
@@ 3 Methodology
We now introduce the SAGMM in detail, with the subsections focusing its key components. Figure 1 illustrates the SAGMM architecture. Input features are first enhanced with structural context, and expert selection is performed by a gating network using Simple Global Attention (SGA) (Wu et al. 2023b) and learnable thresholds. The selected expert outputs are projected into a shared space, weighted by gate scores, and aggregated, while low-importance experts are pruned based on an adaptive threshold $\eta$. The training phase of SAGMM in outlined in Algorithm 1.
@@ 3.1 Pool of Experts
Recent surveys (Chen et al. 2020; Ju et al. 2024) categorize GNNs along three key dimensions: propagation strategy (spectral vs. spatial), aggregation mechanism (mean vs. attention), and training setup (transductive vs. inductive). Guided by these dimensions, we include GNNs like GCN (Kipf and Welling 2016) for its spectral filtering and suitability for homophilic graphs, GAT (Veličković et al. 2017) for attention-based neighbor weighting, and GraphSAGE (Hamilton et al. 2017a) for its inductive capability through neighborhood sampling, along with additional models in the expert pool. Such heterogeneity enables the router to select experts that align with the structural and statistical properties of individual instances.
@@
This design is also motivated by the bias-variance tradeoff (Yang et al. 2020), where over-parameterized models can improve generalization by reducing variance. By incorporating multiple GNNs, our framework benefits from this property. Moreover, in line with the No Free Lunch theorem (Goldblum et al. 2023) which states that no single model performs optimally across all tasks, our approach accounts for variations in network topology, homophily, and heterophily that affects GNN performance. We show empirically that by pooling experts, each with an unique message-passing mechanism, within MoE framework, our approach dynamically adapts to diverse graph structures, overcoming limitations of a single model.
@@ 3.2 Gating Network
Limitations of traditional gating mechanism. Several approaches have been proposed for designing gating networks in MoE architectures (Jordan and Jacobs 1994; Clark et al. 2022; Zhou et al. 2022). A widely adopted sparse gating strategy is the noisy top-$k$ gating mechanism (Shazeer et al. 2017), which applies a softmax over the top-$k$ logits produced by a linear transformation of the input. To encourage expert diversity and prevent load imbalance, this approach injects learnable Gaussian noise into the logits, a modification later adapted for graph data in GMoE model.
@@
    The model’s performance is highly dependent on the choice of $k$, as different datasets exhibit varying optimal values as shown in Figure 2(a) for GMoE-GCN (Wang et al. 2023), highlighting the need for extensive hyperparameter tuning. Moreover, using a fixed $k$ forces all nodes to activate the same number of experts, which may not be needed considering structural and feature characteristics. As shown in Figure 2(b), different nodes activate different number of experts as per their requirement, eventually leading to performance gains.
@@
    Existing methods often rely on simple linear projections of node features to compute gating scores, neglecting structural dependencies within the graph. While DA-MoE (Yao et al. 2024) and GraphMoRE (Guo et al. 2025) attempt to overcome this limitation by leveraging GNN-based gating networks, a fundamental drawback remains: these approaches are still unable to effectively capture global graph structural information, which is vital for informed expert selection.
To address these limitations, for each node, TAAG selects experts based on the local and global structural information while dynamically adjusting the number of activated experts to improve efficiency and robustness. We leverage the SGA block from SGFormer (Wu et al. 2023b) to compute attention scores efficiently. While standard attention mechanisms have a quadratic complexity of $O(n^{2})$, SGA operates in linear time $O(n)$ w.r.t. $n$ nodes, making it scalable for medium to large-scale graphs. To incorporate global structural information, we construct the matrix $X^{(g)}$ using the $p$ smallest eigenvectors of the normalized graph Laplacian $L$ (Eq. (4)), providing positional encodings for each node (Rampášek et al. 2022). We do not use diffusion-based methods (Gasteiger et al. 2019; Chamberlain et al. 2021) as they involve expensive preprocessing, assumes homophily and perform poorly in tasks like link prediction. To further strengthen the structural information, we also add local structural features by aggregating 1-hop and 2-hop (approximate) features (Hamilton et al. 2017a; Wang et al. 2020a). The detailed formulations are provided in Eq. (3) and Eq. (4).
@@
where: $W_{Q},W_{K},W_{V}\in\mathbb{R}^{(d+p)\times N}$ are learnable projection matrices for query, key, and value, respectively. We compute the attention-based gating scores using the Simple Global Attention (SGA) mechanism (Wu et al. 2023b):
@@ 3.3 Adaptive Expert Pruning
Auxiliary Loss Functions. To ensure balanced expert utilization, we incorporate importance loss and diversity loss from prior works (Wang et al. 2023; Guo et al. 2024). These losses prevent over-reliance on specific experts by encouraging more even and orthogonal expert activation patterns.
@@ 4 Experiments
Initial Number of Experts ($N_{0}$). For our primary task of node classification, we construct the expert pool by selecting four and eight widely recognized GNN architectures: GCN, JKNet, GraphCNN, MixHop, GAT, SGC, GIN, and GraphSAGE, all of which are established standards for this application (Hu et al. 2020). In the case of other tasks, the expert pool is composed of four GNN models. The decision to use either eight or four experts is guided by the need to balance architectural diversity with computational feasibility. Notably, this choice is consistent with recent findings in the literature (Yun et al. 2024; He et al. 2025), which indicate that configuring MoE models with four or eight experts enables efficient inference and strong model performance.
@@ 4.1 Performance Evaluation
Node Classification. We evaluate on seven standard datasets ranging from small to large-scale, including four homophilic graphs: ogbn-arxiv, ogbn-proteins, ogbn-products, and ogbn-papers100M from the OGB benchmark (Hu et al. 2020), and three heterophilic graphs: Pokec (Jure 2014), Deezer (Rozemberczki and Sarkar 2020), and YelpChi (Mukherjee et al. 2013). As shown in Table 1, SAGMM outperforms all baselines, achieving notable gains on medium to large-scale datasets such as a 5.90% improvement on ogbn-products and 1.57% on ogbn-papers100M, along with consistent accuracy improvements across the remaining benchmarks. Our analysis reveals that SAGMM consistently prunes weaker experts such as GIN across datasets, demonstrating its ability to adaptively select the most effective experts.
@@ 4.2 Ablation Study
We conduct an ablation study to evaluate the contribution of each component in SAGMM by testing four variants: (i) using identical GNN architectures for all experts (SAGMM (w/o diverse)), (ii) replacing our gating module with noisy Top-$k$ gating (SAGMM (top-$k$ gating)), (iii) using Top-Any gating (Guo et al. 2024) having dynamic expert selection (SAGMM (top-any gating)), and (iv) disabling the adaptive expert pruning mechanism (SAGMM (w/o pruning)). As shown in Table 5, the largest drop is observed when expert diversity is removed, confirming its critical role. Substituting our gating mechanism with top-$k$ or top-any variants impairs expert selection quality, while disabling pruning retains performance in the case of ogbl-collab but always leads to higher memory usage. These results underscore the importance of expert diversity, TAAG gating, and pruning in SAGMM’s effectiveness.